User Tools

Site Tools


This wiki is not maintained! Do not use this when setting up AuScope experiments!

Chair: Tiege

Attended: Mas, Simin, Katie, Patrick, Jamie,

Apologies: Arwin, Bryn, Jonny, Lucia, Jonny


General Issues

  • LBA, how is everyone going?
  • Hobart 26m, holding up?
    • It will be in experiments next week, receivers can not be changed remotely. Make sure S/X is on axis before Brett leaves.
    • Still need to run Windstow script independently. Wind sensor still down.
  • Alternate Mark5 still in use at Yg. still using mk5-2yg.
    • Lucas may be running some additional short notice experiments on weekends, therefore, module logistics may be affected. Keep an eye out for emails from Warren regarding modules you should be using for experiment starts.

Handover Issues

  • Receiver pressure at Hobart 26m?
    • Keep and eye on pressure during hot days, may need to engage cryo pump.
/home/www/auscope/opswiki/data/pages/operations/agenda2019_02_08.txt · Last modified: 2018/02/08 02:22 by Tiege McCarthy